OpenStudy (anonymous):

Landon opened a savings account with a deposit of $3,786.45. Nine years later, the balance in the account is $5,135.29. If the interest rate is 3.4%, how often does the interest compound?

jimthompson5910 (jim_thompson5910):

Hint: use the formula A = P(1+r/n)^(n*t)

jimthompson5910 (jim_thompson5910):

In this case P = 3786.45 r = 0.034 t = 9

jimthompson5910 (jim_thompson5910):

Plug these values into the formula A = P(1+r/n)^(n*t) to get A = P(1+r/n)^(n*t) A = 3786.45(1+0.034/n)^(n*9)

jimthompson5910 (jim_thompson5910):

The goal is to find the value of n which makes A equal to 5135.29

jimthompson5910 (jim_thompson5910):

Common of values of n are n = 1 (annual compounding) n = 2 (semi-annual compounding) n = 4 (quarterly compounding) n = 12 (monthly compounding) n = 365 (daily compounding)

jimthompson5910 (jim_thompson5910):

So plug in each value of n listed above into A = 3786.45(1+0.034/n)^(n*9) ------------------------------------------------------- So for instance, if n = 1, then A = 3786.45(1+0.034/n)^(n*9) A = 3786.45(1+0.034/1)^(1*9) A = 5115.84143581059 A = 5115.84 which is NOT the value of 5135.29, so we must move onto other values of n and check them

jimthompson5910 (jim_thompson5910):

This rules out annual compounding as a possible answer choice

jimthompson5910 (jim_thompson5910):

does all this make sense?

OpenStudy (anonymous):

yes thank you

jimthompson5910 (jim_thompson5910):

ok great, glad it does, yw